S. Lee Enterprises produces two models of lawn tractor: SL1 and SL2. The models are both produced in the company’s Louisville factory and go through identical assembly operations. The difference is in the quality (hence, cost) of the parts used. The following data are available for November, the latest month for which information is available:

SL1 SL2 Total

Number assembled 1,850 2,350 4,200

Materials cost per tractor $ 1,450 $ 1,950

Other costs:

Direct labor $ 1,750,000

Supplies 590,000

Other overhead 3,960,000

Required: S. Lee Enterprises uses operations costing and assigns conversion costs based on the number of units assembled. Compute the cost of each model assembled in November.

of each model assembled in November.

SL1 SL2 Total
Operation Cost
Materials Cost
Total Cost
Number of Units
Unit Cost
Answer #1
SL1 SL2 Total
Operation Cost $2,775,000 $3,525,000 $6,300,000
Materials Cost $2,682,500 $4,582,500 $7,265,000
Total Cost $5,457,500 $8,107,500 $13,565,000
Number of Units 1,850 2,350 4,200
Unit Cost 2,950 3,450 6,400

Working notes:

Total conversion cost ($1,750,000 + $590,000 + $3,960,000) = $6,300,000

Divided by Total units assembled = 4,200

Conversion cost per unit = 1,500

Conversion cost allocated:

SL1 = 1,850 * 1,500 = $2,775,000

SL2 = 2,350 * 1,500 = $3,525,000

Material costs:

SL1 = 1,850 * 1,450 = $2,682,500

SL2 = 2,350 * 1,950 = $4,582,500